# Hesitation and Tolerance in Recommender Systems

### v1

@@ Abstract.
User interactions in recommender systems are inherently complex, often involving behaviors that go beyond simple acceptance or rejection. One particularly common behavior is hesitation, where users deliberate over recommended items, signaling uncertainty. Our large-scale surveys, with 6,644 and 3,864 responses respectively, confirm that hesitation is not only widespread but also has a profound impact on user experiences. When users spend additional time engaging with content they are ultimately uninterested in, this can lead to negative emotions, a phenomenon we term as tolerance. The surveys reveal that such tolerance behaviors often arise after hesitation and can erode trust, satisfaction, and long-term loyalty to the platform. For instance, a click might reflect a need for more information rather than genuine interest, and prolonged exposure to unsuitable content amplifies frustration. This misalignment between user intent and system interpretation introduces noise into recommendation training, resulting in suggestions that increase uncertainty and disengagement. To address these issues, we identified signals indicative of tolerance behavior and analyzed datasets from both e-commerce and short-video platforms. The analysis shows a strong correlation between increased tolerance behavior and decreased user activity. We integrated these insights into the training process of a recommender system for a major short-video platform. Results from four independent online A/B experiments demonstrated significant improvements in user retention, achieved with minimal additional computational costs. These findings underscore the importance of recognizing hesitation as a ubiquitous user behavior and addressing tolerance to enhance satisfaction, build trust, and sustain long-term engagement in recommender systems.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 11ccs: Information systems Recommender systems
@@ 1. Introduction
In recent years, the boundaries of Internet applications have been disappearing. A notable cross-domain effect has emerged: the e-commerce platforms increasingly use short videos to promote products and boost consumer spending (Lei et al., 2021; Guo et al., 2022), while social media and video live-streaming platforms have begun engaging in product sales (Lin et al., 2019; Ladhari et al., 2020; Min Zhang and Luo, 2020; Chen et al., 2019b), which have become major revenue sources. This evolution signifies a profound transformation in traditional business competition strategies. The margins of profit opportunities for these platforms are becoming increasingly blurred, shifting the competition back to the essence of the internet: attracting more users, retaining them longer, and keeping them engaged. Platforms that can achieve this holds a greater competitive advantage and business potential (Du et al., 2019; Legner et al., 2019; Yang et al., 2023; Altman and Jiménez, 2019). Consequently, the emphasis on user retention rates has become a consensus among all practitioners (Jannach and Jugovac, 2019; Gomez-Uribe and Hunt, 2016; Wang et al., 2022).
Recommender systems significantly enhance user satisfaction and retention by providing information or products that fits user’s need (Altman and Jiménez, 2019; Jannach and Jugovac, 2019). Platforms such as TikTok, Kuaishou, and Reels (Zannettou et al., 2024; Chang et al., 2023a; Molem et al., 2024) have garnered millions of users and achieved tremendous commercial success by developing exceptional recommender systems focused on user retention. At the same time, traditional industries such as online retail (Chen et al., 2019a; Leino and Räihä, 2007), news reading (Xiao et al., 2022; Zhang et al., 2023b), and music streaming (Feng et al., 2021; Briand et al., 2021) have also introduced recommendation technologies and adjusted their business models. Consequently, the goal of retaining more users i.e., user retention, and maintaining their activity in recommender systems has received unprecedented attention, with these systems being widely used to enhance user stickiness (Wu et al., 2017; Zou et al., 2019; Cai et al., 2023; Xue et al., 2023b; Silveira et al., 2017; Zhang et al., 2021b; Ding et al., 2023a; Gomez-Uribe and Hunt, 2016).
Conventionally, recommender systems classify user responses to recommended content as either interested or disinterested, treating these responses as positive and negative samples during optimization. Because click behaviors are interpreted as signals of interest, Click-Through Rate (CTR) becomes a crucial metric for assessing recommendation accuracy (Zhang et al., 2023a, 2014; Wang et al., 2015). However, the complexity and diversity of user intentions, and even emotions, complicate the correlation between intentions and the limited set of behavior types (Xia et al., 2021). For instance, a click may not only reflect interest but could also indicate a desire for more information before making a judgment. The final evaluation after clicking might still be disinterest, and the user’s attitude during the session may not clearly fall into either the interested or the disinterested category.
As a result, training recommendation models to interpret clicks primarily as positive signals can introduce uncertainty into the system. This issue is closely tied to the interaction design of the platform, especially regarding the amount and type of information exposed to users. Therefore, it is crucial to understand how this phenomenon affects user experience in recommender systems, and whether such sessions represent an overlooked state of user acceptance, which could directly influence long-term user experience, as reflected in user activity and retention rates.
To the best of our knowledge, there is no in-depth research that has thoroughly investigated this phenomenon. One possible reason for this gap is that academia often lacks corresponding signal verification, while industry, due to commercial competition, tends not to disclose too much on technical indicators. The impact of hesitation remains unclear. To address this gap, we conduct a large-scale user study, combined with dataset analysis and online A/B experiments, to examine the effects of indecision and related emotions on recommender systems.
In this study, we aim to address three research questions (RQs): (1) Does an ambiguous state exist between interest and disinterest? (2) If so, how can this state be accurately identified or captured? (3) If captured, can recommender benefit from such behavioral signals to improve retention rates?
To address RQ1, we conducted two comprehensive user surveys, gathering 6,644 and 3,864 responses, respectively. Our findings confirm that: (i) Users indeed experience a state between interest and disinterest, which we term hesitation; (ii) This hesitation, when associated with unproductive costs and decisions of disinterest, can develop into tolerance; (iii) Tolerance may diminish user enthusiasm for the platform and reduce their activities, then potentially lead to platform abandonment. Based on the processed survey data, we conclude the characteristics of tolerance signals and their application in optimizing recommender systems, which partially answered RQ2. The answer to RQ3 is based on offline analysis of existing datasets of recommender systems from e-commerce and short-video sharing platforms. The analysis concludes that an increase in tolerant behavior decreases user activity in recommender systems. We then deployed online A/B tests on a real-world video-sharing platform with millions of users to verify whether incorporating tolerance signals could reduce platform abandonment. The online results show that even low-cost strategies targeting these signals can significantly enhance next-day user retention. Training models to incorporate tolerance signals not only boosts user activity but also improves retention rates, addressing key challenges in industrial recommender systems in a cost-effective manner.
@@ 2. Related Work
Our research is related to user behavior modeling and optimization objectives in recommender systems. User behaviors, such as clicks, views, and purchases, serve as essential data for the recommender system to predict user preferences and satisfaction. Leveraging user behaviors, various optimization objectives have been proposed, i.e., predictions of click-through rate, click conversion rate, dwell time, and even long-term outcomes such as user retention. In this section, we briefly discuss existing research related to user behavior modeling and optimization objectives in recommender systems.
2.1. User Behaviors Modelling in Recommender System
Among various ways to implement and refine recommender systems, user behavior is crucial for training and evaluating recommendation models (Wu et al., 2022; Chen et al., 2023; Wang et al., 2021a). Both academia and industry concur on the key behaviors that signal interest or disinterest (Wang et al., 2023; Frolov and Oseledets, 2016; Park and Lee, 2022; Zhao et al., 2018; Liu et al., 2010). However, there remains a notable gap in the study of more nuanced, long-tail behaviors, with even less attention given to the deeper emotional interpretations of user actions (Tkalčič, 2018; Xia et al., 2021).
Preliminary work by Tkalčič (2018) provides a theoretical overview and practical insights into the impact of emotions and personality on recommender systems, concluding that user emotions significantly influence the generation and optimization of recommendations. Recently, a few studies (Pan et al., 2023; Zannettou et al., 2024; Li et al., 2022; Curmei et al., 2022) delve into behavioral interpretation and modeling through exploring the relationships among user interests (Zannettou et al., 2024), activities (Pan et al., 2023; Li et al., 2022), dwell time (Zhan et al., 2022; Zhao et al., 2024), etc., and developed various optimization objectives (Curmei et al., 2022; Li et al., 2022). Specifically, some studies propose to optimize user behavior modeling by focusing on mitigating the systematic and algorithmic biases in recommender systems (Chen et al., 2023). The systematic biases are usually suppressed by promoting equity of attention for items or improving model performances for subgroups (Beutel et al., 2019; Biega et al., 2018; Mehrotra et al., 2018; Zehlike and Castillo, 2020). The algorithm biases are commonly solved from the perspective of causal reasoning (Pearl, 2012). These solutions are roughly categorized into causal embedding (Bonner and Vasile, 2018; Guo et al., 2019; Zhang et al., 2023b), inverse propensity weighting (Christakopoulou et al., 2020; Kool et al., 2020; Bietti et al., 2021), and causal intervention (Wang et al., 2021b; Yang et al., 2021; Zhang et al., 2021a) based methods.
In addition, there are studies exploring to mine more useful user features (Cho et al., 2019; Park et al., 2017; Xiao et al., 2022) or devise multi-task frameworks (Chang et al., 2023b; Ma et al., 2018b; Tang et al., 2020; Zhao et al., 2019) to improve the personalization of recommendations and enhance user satisfaction.
Despite the valuable insights these studies provide, they often concentrate on specific optimization objectives, leaving a gap in understanding user behavioral dynamics throughout the decision-making process. To address this, our research is the first to introduce and formally define two critical concepts, i.e., hesitation and tolerance signals, within the context of recommender systems. These signals are applied in practice to capture subtle user behaviors that have been largely overlooked. Our work fills a gap in both academic and industry research by offering foundational insights into user decision-making patterns that have not yet been systematically explored.
2.2. Optimization Objective in Recommender System
@@
Click is one of the key user behaviors in recommender system. The conversion of a click can lead to an item purchase or an advertisement view on an online platform, ultimately generating business revenue. Accurately predicting whether a user will click on an item in a given context is crucial for optimizing how items are displayed to maximize business outcomes (Zhou et al., 2018, 2019; Lin et al., 2023). This task, known as click-through rate (CTR) prediction, has been extensively studied, with numerous models proposed over the years.
@@
Conventional CTR models, such as Factorization Machines (FM) (Rendle, 2010), model pairwise feature interactions for binary classification of clicks and non-clicks. Other approaches (Guo et al., 2017; Lian et al., 2018) extend these models by incorporating higher-order feature interactions through the use of deep neural networks (DNNs). Zhou et al. (2018) further enhance CTR models by introducing a local activation unit to capture the relevance between candidate ads and users’ historical behaviors. In DIEN (Zhou et al., 2019), the authors propose leveraging auxiliary loss in CTR models to supervise the learning of user behavior sequences at each time step. Note that, traditional CTR prediction methods typically treat all clicks equally, without distinguishing between the quality of clicks. Nevertheless, studies (Xie et al., 2023; Wen et al., 2019) also highlight that while clicks reflect users’ initial impressions, they do not necessarily indicate user satisfaction with items after further exploration. These studies suggest that post-click behaviors should be considered in CTR models to better represent user preferences.
@@
In addition to predicting click-through rate (CTR), click conversion rate (CVR) prediction is also considered a crucial task in recommender systems. Generally, click conversion relates to the “click-to-purchase” path, which directly impacts final revenue. However, predicting CVR is fundamentally more challenging than predicting CTR due to issues such as selection bias and the sparsity of positive samples in the data. To address these challenges, Ma et al. (2018a) propose a multi-task learning approach that simultaneously trains CTR and CVR models, allowing the CVR prediction task to benefit from the abundant positive samples available in the CTR prediction task. Similarly, the authors of (Wen et al., 2020) introduce a multi-task model based on conditional probability rules to enhance CVR learning. Wen et al. (2021) further propose to decompose user behaviors into detailed actions, such as clicking on item images and engaging in conversations with sellers, providing richer signals for CVR prediction.
@@
When it comes to video recommendation scenario, beyond clicks, watch-time prediction is regarded as a more representative metric of user engagement (Wu et al., 2018; de Sá et al., 2021; Covington et al., 2016). Covington et al. (2016) presents an industry-standard solution to watch-time prediction, where the regression problem is transformed into the weighted logistic regression. Zhan et al. (2022) further analyzes the duration bias in watch-time prediction via a causal graph and duration-deconfounded quantile-based framework. Zhao et al. (2024) introduces the counterfactual watch time concept to measure the potential duration bias and devise a corresponding counterfactual watch model to solve this problem.

### v2

@@ Abstract.
Users’ interactions with recommender systems often involve more than simple acceptance or rejection. We highlight two overlooked states: hesitation, when people deliberate without certainty, and tolerance, when this hesitation escalates into unwanted engagement before ending in disinterest. Across two large-scale surveys ($N=6,644$ and $N=3,864$), hesitation was nearly universal, and tolerance emerged as a recurring source of wasted time, frustration, and diminished trust. Analyses of e-commerce and short-video platforms confirm that tolerance behaviors, such as clicking without purchase or shallow viewing, correlate with decreased activity. Finally, an online field study at scale shows that even lightweight strategies treating tolerance as distinct from interest can improve retention while reducing wasted effort. By surfacing hesitation and tolerance as consequential states, this work reframes how recommender systems should interpret feedback, moving beyond clicks and dwell time toward designs that respect user value, reduce hidden costs, and sustain engagement.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791865isbn: 979-8-4007-2278-3/2026/04ccs: Information systems Recommender systems
@@ 1. Introduction
Imagine opening a recommended video that seems appealing at first. You pause, watch for a while, and then realize it is not what you wanted. The time and attention you spent are gone, an experience of disappointment rather than value. Such moments are common, yet most recommender systems (RecSys) log them simply as a “click,” treating them as success.
Over the past decade, platforms from e-commerce (Chen et al., 2019; Leino and Räihä, 2007) to short-video streaming (Zannettou et al., 2024; Chang et al., 2023a; Molem et al., 2024) have invested heavily in recommender systems to increase retention and engagement (Du et al., 2019; Legner et al., 2019; Yang et al., 2023; Altman and Jiménez, 2019; Jannach and Jugovac, 2019). While these systems are celebrated for their reach and profitability, the signals that drive them remain strikingly coarse. For example, clicks are routinely collapsed into binary categories of positive (interested) or negative (disinterested) in recommender systems, erasing the lived nuances of curiosity, doubt, and regret before “click”. As a result, metrics like Click-Through Rate (CTR) have become the dominant proxy for user satisfaction (Zhang et al., 2023c, 2014; Wang et al., 2015), even though they frequently capture hesitation or frustration rather than genuine preference.
The complexity and diversity of user intentions, and even emotions, complicate the correlation between intentions and the limited set of behavior types (Xia et al., 2021), which is deeply problematic. Users do not simply react to algorithms; they strategically explore, second-guess, and occasionally persist with content they do not truly want. These moments are costly (not in computation, but in human terms): time wasted, trust eroded, agency undermined. Yet they remain invisible in current optimization frameworks.
Addressing this gap requires shifting attention from model architectures to the dynamics of user-recommender interaction. Decades of HCI research demonstrate that users act strategically, form folk theories of algorithms, and actively probe or steer them (Eslami et al., 2016; Rader and Gray, 2015; DeVito et al., 2017). Perspectives on agency, trust, and cognitive load further caution that engagement metrics may signal doubt, frustration, or compulsion rather than satisfaction (Lukoff et al., 2021; Pirolli, 2007). These insights suggest that current RecSys models overlook critical interaction states between interest and disinterest, such as moments of hesitation, second-guessing, and persistence despite dissatisfaction.
In this paper, we investigate two such states. Hesitation arises when users deliberate over recommended content, clicking without certainty. Tolerance occurs when hesitation escalates into unwanted engagement, leading to frustration and wasted effort. These states are consequential because they introduce hidden interaction costs, time, cognitive effort, and negative emotions, that erode trust and retention while remaining invisible to current optimization metrics. By examining these states, we aim to broaden how recommender systems interpret feedback, shifting attention from maximizing clicks to minimizing wasted effort. To investigate these dynamics, we pose three research questions: RQ1: Does an intermediate state exist between interest and disinterest in recommender interactions? RQ2: If so, how can this state be systematically identified and modeled? RQ3: If captured, can recommender systems leverage these signals to improve user retention?
We conducted a multi-method, multi-scale study comprising (1) large-scale surveys, (2) behavioral log analyses, and (3) a field deployment. For RQ1, we designed two surveys ($N=6,644$ and $N=3,864$). Findings show that: (i) users indeed experience an ambiguous state, which we term hesitation; (ii) hesitation, when coupled with unproductive costs and eventual disinterest, often develops into tolerance; and (iii) tolerance undermines enthusiasm, reduces activity, and may ultimately lead to abandonment. From these survey findings, we characterized tolerance signals and demonstrated their potential for recommender optimization, partially addressing RQ2.
To further address RQ2 and RQ3, we analyzed behavioral datasets from e-commerce and short-video platforms. Results indicate that higher tolerance correlates with reduced activity. Finally, we validated our approach through large-scale online A/B testing on a video-sharing platform with millions of users. The results demonstrate that even lightweight strategies, such as relabeling tolerance signals, can produce measurable improvements in next-day retention. Training models with these signals not only boosts activity but also enhances retention, offering a cost-effective response to long-standing industrial challenges.
Across multiple deployments, these adjustments yielded consistent gains, with next-day retention increasing by up to 0.67%, a figure that, while modest, represents a substantial improvement at industrial scale. Beyond technical optimization, our work reframes hesitation and tolerance as consequential interaction states. By doing so, we move recommender system research toward a more human-centered perspective, highlighting how systems can respect users’ time, preserve trust, and foster sustainable engagement.
Our contributions are fourfold: (1) we conceptualize and operationalize hesitation and tolerance as measurable intermediate interaction states, empirically establishing their prevalence and direct impact on retention; (2) we provide converging evidence, across surveys, behavioral logs, and an industrial-scale A/B test, that these states depress trust, satisfaction, and engagement; (3) we propose and validate modeling strategies that incorporate tolerance signals as weak positives/negatives, bridging RecSys optimization with user-centered concerns about user cost and agency; and (4) we demonstrate practical impact: modest relabeling of tolerance-related signals both improves retention and reduces wasted effort, offering a pathway to design recommenders that respect user time and foster sustainable engagement.
@@ 2. Related Work
Research in user behavior modeling and optimization objectives in recommender systems exist in both RecSys and SIGCHI communities. User behaviors, such as clicks, views, and purchases, serve as essential data for the recommender system to predict user preferences and satisfaction. Leveraging user behaviors, various optimization objectives have been proposed, i.e., predictions of click-through rate, click conversion rate, dwell time, and even long-term outcomes such as user retention. In the meantime, prior work have examined how users understand, experience, and adapt to recommender systems, focusing less on algorithmic optimization and more on users’ mental models, agency, and the social and psychological consequences of interacting with algorithmic recommendations. In this section, we discuss existing work related to user behavior modeling in both RecSys and SIGCHI research, as well as optimization objectives in recommender systems (RecSys)
2.1. Modeling User Behaviors in RecSys
Early recommender systems research often emphasized how users form and act on their mental models of algorithms. Studies have shown that people rarely treat recommendations as neutral suggestions; instead, they construct “folk theories” of how algorithms operate and attempt to steer or correct them through their actions. For example, users may deliberately click, like, or hide content to signal preferences, even when these actions do not align with their actual interest, reflecting a belief that such signals will shape future recommendations (Kulesza et al., 2012; Eslami et al., 2016; Rader and Gray, 2015; DeVito et al., 2017). This line of work highlights that user behaviors in recommender systems cannot be interpreted as direct expressions of preference, but are deeply entangled with users’ understandings, expectations, and strategies for influencing algorithms. Beyond conceptual models, prior work has also examined hesitation as a micro-level social signal manifested through nonverbal cues such as hand movements and gaze dynamics in conversational recommendation settings (Vodlan et al., 2015).
Building on this perspective, subsequent user-centered studies have examined how people decide whether to accept or reject recommendations, especially in contexts like news feeds, video streaming, or e-commerce. Research shows that users often deliberate before acting on a recommendation, weighing factors such as trust in the platform, perceived relevance, and anticipated consequences of engagement (Wang et al., 2022; Lukoff et al., 2021). When users do click, these behaviors are not always equivalent to satisfaction or endorsement; instead, actions like briefly opening a video, skipping ahead, or abandoning a purchase may signal ambivalence or doubt (Chaudhary et al., 2022; Zannettou et al., 2024). Work in this area has drawn on broader user-centered theories of agency, cognitive load, and information foraging to explain why clicks and views can mask underlying uncertainty or disengagement (Pirolli, 2007; Card et al., 2001).
Complementing these user-centered perspectives, recommender systems research has long treated user behavior as the primary input for training and evaluating models. User clicks, views, and purchases serve as essential signals for predicting preferences and satisfaction, and a large body of work has refined optimization objectives around metrics such as click-through rate, conversion rate, dwell time, and retention (Wu et al., 2022; Chen et al., 2023; Wang et al., 2021a, 2023a; Frolov and Oseledets, 2016; Park and Lee, 2022; Zhao et al., 2018; Liu et al., 2010). While these behavioral proxies have proven useful, they often treat user actions as direct indicators of interest, overlooking the more nuanced and long-tail signals that may reflect ambivalence, disengagement, or emotional states (Tkalčič, 2018; Xia et al., 2021). For instance, preliminary work has explored the impact of emotions and personality on recommender systems (Tkalčič, 2018), while more recent efforts have examined how user interests (Zannettou et al., 2024), activities (Pan et al., 2023; Li et al., 2022), and temporal signals such as dwell time (Zhan et al., 2022; Zhao et al., 2024) shape system performance and evaluation.
Another stream of research has focused on mitigating systematic and algorithmic biases in user behavior modeling. Systematic biases are often addressed by promoting equity of attention for items or improving model performance across subgroups (Beutel et al., 2019; Biega et al., 2018; Mehrotra et al., 2018; Zehlike and Castillo, 2020), while algorithmic biases are increasingly approached from the perspective of causal reasoning (Pearl, 2012). This has led to the development of causal embedding (Bonner and Vasile, 2018; Guo et al., 2019; Zhang et al., 2023a), inverse propensity weighting (Christakopoulou et al., 2020; Kool et al., 2020; Bietti et al., 2021), and causal intervention (Wang et al., 2021b; Yang et al., 2021; Zhang et al., 2021a) methods to correct confounding factors and improve robustness. Beyond bias mitigation, researchers have also explored mining richer user features (Cho et al., 2019; Park et al., 2017; Xiao et al., 2022) and designing multi-task frameworks (Chang et al., 2023b; Ma et al., 2018a; Tang et al., 2020; Zhao et al., 2019) to improve personalization and enhance user satisfaction.
Despite these advances, both HCI and RecSys research tend to focus on either users’ conceptual models of algorithms or measurable optimization objectives tied to observable behaviors. Less attention has been given to the dynamic decision-making processes that precede or underlie these behaviors. In particular, prior work has not systematically addressed the dynamics of hesitation, when users deliberate at the threshold of action, or tolerance, when they continue engaging with content they do not truly want. Our research introduces and formalizes these two critical concepts, filling a gap in both academic and industry literature by offering foundational insights into subtle user decision-making patterns that existing models and theories have largely overlooked.
2.2. Optimization Objectives in RecSys
Optimizing recommender systems is a complex endeavor that necessitates considering multiple facets and perspectives, and it is accompanied by trade-offs between evaluation metrics. A comprehensive evaluation must address both system-centric aspects and user-centric aspects. Furthermore, a strong focus, particularly in industry, is placed on business effectiveness and overall goals, such as increasing revenue and customer engagement (Zangerle and Bauer, 2022; Zou and Sun, 2025).
@@ 2.2. Optimization Objectives in RecSys
Click is one of the key user behaviors in recommender systems. The conversion of a click can lead to an item purchase or an advertisement view on an online platform, ultimately generating business revenue. Accurately predicting whether a user will click on an item in a given context is crucial for optimizing how items are displayed to maximize business outcomes (Zhou et al., 2018, 2019; Lin et al., 2023). This task, known as click-through rate (CTR) prediction, has been extensively studied, with numerous models proposed over the years.
@@
Conventional CTR models, such as Factorization Machines (FM; (Rendle, 2010)), model pairwise feature interactions for binary classification of clicks and non-clicks. Other approaches (Guo et al., 2017; Lian et al., 2018) extend these models by incorporating higher-order feature interactions through the use of deep neural networks (DNNs). Zhou et al. (2018) further enhance CTR models by introducing a local activation unit to capture the relevance between candidate ads and users’ historical behaviors. In DIEN (Zhou et al., 2019), the authors propose leveraging auxiliary loss in CTR models to supervise the learning of user behavior sequences at each time step.
However, traditional CTR prediction methods typically treat all clicks equally, without distinguishing between the quality of clicks. Studies have shown that while clicks reflect users’ initial impressions, they do not necessarily indicate user satisfaction with items after further exploration (Xie et al., 2023; Wen et al., 2019). In HCI research, this distinction aligns with findings that actions such as clicking, skipping, or briefly viewing may stem from ambivalence or testing the system rather than genuine interest (Lukoff et al., 2021; Chaudhary et al., 2022). Together, these insights suggest that post-click behaviors should be considered in CTR models to better capture the nuance of user preferences.
@@
Beyond CTR, click conversion rate (CVR) prediction is also a crucial task in recommender systems. Generally, click conversion relates to the “click-to-purchase” path, which directly impacts final revenue. However, predicting CVR is fundamentally more challenging than predicting CTR due to issues such as selection bias and the sparsity of positive samples in the data. To address these challenges, Ma et al. (2018b) propose a multi-task learning approach that simultaneously trains CTR and CVR models, allowing the CVR prediction task to benefit from the abundant positive samples available in the CTR prediction task. Similarly, the authors of (Wen et al., 2020) introduce a multi-task model based on conditional probability rules to enhance CVR learning. Wen et al. (2021) further propose to decompose user behaviors into detailed actions, such as clicking on item images and engaging in conversations with sellers, providing richer signals for CVR prediction. Such decompositions echo work on understanding the micro-decisions that shape users’ mental models of systems and influence their willingness to trust or act on recommendations (Rader and Gray, 2015; Eslami et al., 2016).
@@
When it comes to video recommendation scenarios, beyond clicks, watch-time prediction is regarded as a more representative metric of user engagement (Wu et al., 2018; de Sá et al., 2021; Covington et al., 2016). Covington et al. (2016) presents an industry-standard solution to watch-time prediction, where the regression problem is transformed into the weighted logistic regression. Zhan et al. (2022) further analyze the duration bias in watch-time prediction via a causal graph and a duration-deconfounded quantile-based framework. Zhao et al. (2024) introduce the counterfactual watch time concept to measure the potential duration bias and devise a corresponding counterfactual watch model to solve this problem. Prior work has questioned whether watch time or dwell time truly equate to satisfaction, pointing to cases where prolonged engagement may actually reflect frustration, compulsion, or lack of alternatives (Lukoff et al., 2021; Pirolli, 2007).
In addition to click-based objectives, a wide range of interaction signals are used in both academic and industrial recommender systems. Long-CTR, which counts only clicks that lead to sustained consumption, has long been deployed in video and content platforms as a more robust engagement signal than raw clicks (Davidson et al., 2010; Jadon and Patil, 2024). Beyond this, completion-based measures—such as watched-percentage or scrolled-percentage ratios—are widely adopted using distributional labeling or causal debiasing techniques (Zhan et al., 2022; Zhang et al., 2023b). Low-engagement indicators, including skip rate and early exits, are frequently leveraged to capture disengagement in short-video recommenders (Gong et al., 2022). At the sequence level, session-duration dynamics and next-item transitions are commonly modeled through long-sequence or session-based architectures (Si et al., 2024; Wu et al., 2024; Mei et al., 2022). Platforms also rely on explicit feedback such as likes, saves, and ratings to refine user preferences (Qin et al., 2021; Shao et al., 2024). Finally, novelty and serendipity-oriented objectives are used to encourage exploratory consumption and mitigate fatigue (Li et al., 2020; Lin et al., 2022; Wang et al., 2023b). These metrics collectively illustrate the breadth of signals used to assess engagement and value in practice, and they highlight the limitations of interpreting surface-level behaviors as direct indicators of interest—motivating the need to examine interaction-level states such as hesitation and tolerance.